# 7-3-1 Hongo, Bunkyo-ku, Tokyo 113-0033, Japan

### v3

@@ 2.2 Universal results
In this subsection, we show detailed examples of calculating the modular $S$ matrices, fusion rules, and the disk two-point functions using with Cardy boundary states, using free-field approaches. We will use the simplest solution to the neutrality conditions for our calculation. We only show the integral contributions from a single Ishibashi state, without writing the the final summation of each contribution.
@@ 2.3.1 Virasoro minimal models
It is easy to verify that $S^{2}=C=I$. The non-obvious non-zero fusion rules are
@@
Consider primary disk two-point functions $\langle b_{(r,s)}|\Phi_{(2,1)}(z_{1},\widebar{z}_{1})\Phi_{(2,1)}(z_{2},\widebar{z}_{2})|0\rangle$. It consists of two types contributions: $I_{(2,1)(2,1)}^{(1,1)}:=\langle\langle L_{(1,1)}|\Phi_{(2,1)}(z_{1},\widebar{z}_{1})\Phi_{(2,1)}(z_{2},\widebar{z}_{2})|0\rangle$ and $I_{(2,1)(2,1)}^{(2,1)}:=\langle\langle L_{(2,1)}|\Phi_{(2,1)}(z_{1},\widebar{z}_{1})\Phi_{(2,1)}(z_{2},\widebar{z}_{2})|0\rangle$. A free-field vertex operator realization of the contribution $I_{(2,1)(2,1)}^{(1,1)}$ is
@@
It can be easily verified that $S^{2}=C=I$. The non-obvious non-zero fusion rules are
@@ 2.3.2 $\mathcal{W}_{3}(p,p^{\prime})$ minimal models
It is easy to verify that $S^{2}=C=I$. The non-obvious non-zero fusion rules are given by
@@
The simplest unitary $\mathcal{W}_{3}$ minimal model is the $\mathcal{W}_{3}(5,4)$ minimal model. The diagonal $\mathcal{W}_{3}(5,4)$ minimal model contains a non-diagonal Virasoro sub-theory, which is the three-state Potts model. Studying the three-state Potts model using free-field approaches to $\mathcal{W}_{3}$ minimal models has been given in Caldeira:2003zz . We consider the free-field approach to disk two-point functions of the diagonal and the charge-conjugated $\mathcal{W}_{3}(5,4)$ minimal models. The $\mathcal{W}_{3}(5,4)$ Kac-table is
@@ 2.3.3 The $\mathcal{W}_{4}(7,4)$ minimal model
$\sum_{k,l,m}|z_{1}|^{\frac{8}{7}+2m_{2}}(1-|z_{1}|^{2})^{\frac{4}{7}}\frac{\Gamma(\frac{1}{7})}{\Gamma(\frac{8}{7})}\frac{(\frac{1}{7})_{\sum k_{i}}(\frac{4}{7})_{k_{1}}(\frac{4}{7})_{k_{2}}}{k_{1}!k_{2}!\frac{8}{7})_{\sum k_{i}}}\frac{\Gamma(a^{\prime})\Gamma(\frac{3}{7})}{\Gamma(c^{\prime})}\frac{(a^{\prime})_{l}(\frac{4}{7})_{l}}{l!(c^{\prime})_{l}}\frac{\Gamma(a)\Gamma(\frac{15}{7})}{\Gamma(c)}$
@@
$\int_{0}^{1}dt\>t^{-\frac{8}{7}}(1-t)^{0}(1-tz^{(1)}\widebar{z}_{1})^{-\frac{4}{7}}.$ (108)
@@ 2.3.4 The $\mathcal{W}\big{[}\widehat{G}_{2}\big{]}(6,7)$ minimal model
The simple roots and coroots are $\alpha_{1}=\alpha_{1}^{\vee}=[2,-3]$ and $\alpha_{2}=\frac{1}{3}\alpha_{2}^{\vee}=[-1,2]$. Hence, $r^{\vee}=3$, and the $G_{2}$ Weyl vector and dual Weyl vector are $\rho=[1,1]$ and $\rho^{\vee}=[1,3]$. The highest root $\theta=2\alpha_{1}+3\alpha_{2}=2\alpha_{1}+\alpha_{2}^{\vee}$. The Coxeter and the dual of Coxeter number of $\widehat{G}_{2}$ are $h=6$ and $h^{\vee}=4$, respectively. $|P/Q^{\vee}|=\det(\alpha_{i}^{\vee},\alpha_{j}^{\vee})=3$. The $G_{2}$ Weyl group has a total of 12 elements, generated by the two simple Weyl reflections $w_{1}=\begin{pmatrix}-1&0\\ 3&1\end{pmatrix}$ and $w_{2}=\begin{pmatrix}1&1\\ 0&-1\end{pmatrix}$. The longest Weyl element has length $l(w_{l})=6$. The zeroth Dynkin label of a $\widehat{G}_{2}$ weight vector $\widehat{\lambda}$ is given by
@@
$\widehat{\Lambda}^{(+i)}\in\widehat{P}_{+}^{p^{i}-h^{\vee i}},\quad\widehat{\Lambda}^{(-i)}\in\widehat{P}_{+}^{\vee\>(p^{\prime})^{i}-h^{i}}.$ (124)
@@ Appendix B A Free-field approach to coset $\mathcal{W}$ minimal Ishibashi states
$L_{\widehat{\Lambda}}^{\widehat{g}}=\bigoplus_{\widehat{\lambda}}\big{(}L_{(\widehat{\Lambda},\widehat{\lambda})}^{\widehat{g}/\widehat{h}}\otimes L_{\widehat{\lambda}}^{\widehat{h}}\big{)}$, (130)
@@
$|F_{\widehat{w}^{(i)}\ast\widehat{\Lambda}^{(i)}}\rangle\rangle_{\Omega}=\exp{\Big{[}\sum_{n\in\mathbb{Z}_{+}}\frac{-1}{n}a_{-n}\cdot V_{\Omega}\cdot\widebar{a}_{-n}\Big{]}}|p\otimes\widebar{p}\rangle,\quad p=\alpha_{+}\widehat{w}^{(i)}\ast\widehat{\Lambda}^{(i)}.$ (139)
@@
$|F^{\beta^{\alpha}\gamma^{\alpha}}\rangle\rangle_{I}=\exp{\Big{[}-\sum_{\alpha\in\Delta_{+}}\sum_{n\in\mathbb{Z}_{+}}(\beta_{-n}\widebar{\gamma}_{-n}+\gamma_{-n}\widebar{\beta}_{-n})\Big{]}}|0\otimes\widebar{0}\rangle.$ (140)

### v4

@@ 2.2 Universal results
In this subsection, we show detailed examples of calculating the modular $S$ matrices, fusion rules, and the disk two-point functions using with Cardy boundary states, using free-field approaches. We will use the simplest solution to the neutrality conditions as our integrand. We only write down the integral results of each Ishibashi states, and omit the final summation of all contribution step.
@@ 2.3.1 Virasoro minimal models
They both have the expected Virasoro minimal primary conformal weights
$h_{(r,s)}=\frac{(rp^{\prime}-sp)^{2}-(p-p^{\prime})^{2}}{4pp^{\prime}}.$ (53)
@@
It is easy to verify that $S^{2}=C=I$. The non-zero fusion rules are
@@
Consider primary disk two-point functions $\langle b_{(r,s)}|\Phi_{(2,1)}(z_{1},\widebar{z}_{1})\Phi_{(2,1)}(z_{2},\widebar{z}_{2})|0\rangle$. It consists of two types contributions: $I_{(2,1)(2,1)}^{(1,1)}=\langle\langle L_{(1,1)}|\Phi_{(2,1)}(z_{1},\widebar{z}_{1})\Phi_{(2,1)}(z_{2},\widebar{z}_{2})|0\rangle$ and $I_{(2,1)(2,1)}^{(2,1)}=\langle\langle L_{(2,1)}|\Phi_{(2,1)}(z_{1},\widebar{z}_{1})\Phi_{(2,1)}(z_{2},\widebar{z}_{2})|0\rangle$. A free-field vertex operator realization of the contribution $I_{(2,1)(2,1)}^{(1,1)}$ is
@@
It can be easily verified that $S^{2}=C=I$. The non-zero fusion rules are
@@ 2.3.2 $\mathcal{W}_{3}(p,p^{\prime})$ minimal models
It is easy to verify that $S^{2}=C=I$. The non-obvious non-zero fusion rules included in
@@
The simplest unitary $\mathcal{W}_{3}$ minimal model is the $\mathcal{W}_{3}(5,4)$ minimal model. The diagonal $\mathcal{W}_{3}(5,4)$ minimal model contains a non-diagonal Virasoro sub-theory, which is the three-state Potts model. Studying the boundary three-state Potts model using free-field approaches to $\mathcal{W}_{3}$ minimal models has been given in Caldeira:2003zz . We consider the free-field approach to disk two-point functions of the diagonal and the charge-conjugated $\mathcal{W}_{3}(5,4)$ minimal models. The $\mathcal{W}_{3}(5,4)$ Kac-table is
@@ 2.3.3 The $\mathcal{W}_{4}(7,4)$ minimal model
$\sum_{k,l,m}|z_{1}|^{\frac{8}{7}+2m_{2}}(1-|z_{1}|^{2})^{\frac{4}{7}}\frac{\Gamma(\frac{1}{7})}{\Gamma(\frac{8}{7})}\frac{(\frac{1}{7})_{\sum k_{i}}(\frac{4}{7})_{k_{1}}(\frac{4}{7})_{k_{2}}}{k_{1}!k_{2}!(\frac{8}{7})_{\sum k_{i}}}\frac{\Gamma(a^{\prime})\Gamma(\frac{3}{7})}{\Gamma(c^{\prime})}\frac{(a^{\prime})_{l}(\frac{4}{7})_{l}}{l!(c^{\prime})_{l}}\frac{\Gamma(a)\Gamma(\frac{15}{7})}{\Gamma(c)}$
@@
$\int_{0}^{1}dt\>t^{-\frac{8}{7}}(1-tz^{(1)}\widebar{z}_{1})^{-\frac{4}{7}}.$ (108)
@@ 2.3.4 The $\mathcal{W}\big[\widehat{G}_{2}\big](6,7)$ minimal model
The simple roots and coroots are $\alpha_{1}=\alpha_{1}^{\vee}=[2,-3]$ and $\alpha_{2}=\frac{1}{3}\alpha_{2}^{\vee}=[-1,2]$. Hence, $r^{\vee}=3$, and the $G_{2}$ Weyl vector and dual Weyl vector are $\rho=[1,1]$ and $\rho^{\vee}=[1,3]$. The highest root $\theta=2\alpha_{1}+3\alpha_{2}=2\alpha_{1}^{\vee}+\alpha_{2}^{\vee}$. The Coxeter and the dual of Coxeter number of $\widehat{G}_{2}$ are $h=6$ and $h^{\vee}=4$, respectively. $|P/Q^{\vee}|=\det(\alpha_{i}^{\vee},\alpha_{j}^{\vee})=3$. The $G_{2}$ Weyl group has a total of 12 elements, generated by the two simple Weyl reflections $w_{1}=\begin{pmatrix}-1&0\\ 3&1\end{pmatrix}$ and $w_{2}=\begin{pmatrix}1&1\\ 0&-1\end{pmatrix}$. The longest Weyl element has length $l(w_{l})=6$. The zeroth Dynkin label of a $\widehat{G}_{2}$ weight vector $\widehat{\lambda}$ is given by
@@
$\widehat{\Lambda}^{(+)i}\in\widehat{P}_{+}^{p^{i}-h^{\vee i}},\quad\widehat{\Lambda}^{(-)i}\in\widehat{P}_{+}^{\vee\>(p^{\prime})^{i}-h^{i}}.$ (124)
@@ Appendix B A Free-field approach to coset $\mathcal{W}$ minimal Ishibashi states
[…]
@@
$|F_{\widehat{w}^{(i)}\ast\widehat{\Lambda}^{(i)}}\rangle\rangle_{\Omega}=\exp{\Big[\sum_{n\in\mathbb{Z}_{+}}\frac{-1}{n}a_{-n}\cdot V_{\Omega}\cdot\widebar{a}_{-n}\Big]}|p\otimes\widebar{p}\rangle.$ (139)
@@
$|F^{\beta^{\alpha}\gamma^{\alpha}}\rangle\rangle_{I}=\exp{\Big[-\sum_{\alpha\in\Delta_{+}}\sum_{n\in\mathbb{Z}_{+}}(\beta_{-n}\widebar{\gamma}_{-n}+\gamma_{-n}\widebar{\beta}_{-n})\Big]}|0\otimes\widebar{0}\rangle,$ (140)
$|F^{\beta^{\alpha}\gamma^{\alpha}}\rangle\rangle_{U}=\exp{\Big[\sum_{\alpha\in\Delta_{+}}\sum_{n\in\mathbb{Z}_{+}}(\beta_{-n}\widebar{\beta}_{-n}-\gamma_{-n}\widebar{\gamma}_{-n})\Big]}|0\otimes\widebar{0}\rangle.$ (141)